A wire of length 7 π‘š and radius 2.37 π‘šπ‘š is stretched by a load of 12370 π‘”π‘š and increased by 1.78% of the original length. Find the stress, strain, and Young’s modulus of the material of the wire.Β 


1
Expert's answer
2021-12-06T09:47:45-0500

Determine the required values: stress "\sigma", strain "\epsilon", and Young's modulus "Y":


"\sigma=\frac{F}{A}=\frac {mg}{\pi r^2}=6.87\text{ MPa}.\\\space\\
\epsilon=1.78/100=0.0178.\\\space\\
E=\frac{\sigma}{\epsilon}=386\text{ MPa}."